OpenStudy (anonymous):

. What is the fractional equivalent of 0.6?

OpenStudy (anonymous):

0.6=6/10=3/5

OpenStudy (anonymous):

.6 = 6/10 = 3/5

OpenStudy (saifoo.khan):

3/5

OpenStudy (anonymous):

Ah.. ninja'd by mrjoselson :D

OpenStudy (anonymous):

I'll copy your answer next time :)